The High Energy Frontier

CDF II at the Tevatron Collider

Pekka Sinervo University of Toronto

Overview

• Introduction

• Tevatron and CDF upgrades • Top Quark Physics

• B Physics Goals

• Higgs Search

• Summary

The Standard Model

The ingredients of the SM are

• 12 fermions – 6 quarks (u, d, s, c, b, t)

– 6 leptons (e^{−}, ν_{e}, μ, ν_{μ}, τ, ν_{τ})

• 3 electroweak bosons (W^{+} , W^{−}, Z^{∘}, γ)

• 8 gluons to describe the strong force

• ?? to describe gravity

Enormously successful paradigm since ∼ 1975

• Predicted rich phenomenology

• Accurate! (eg. δ_{M_{z}}/M_{z} ∼ 8 × 10^{−5})

• Successfully predicted existence of charm and top quark

Higgs boson the only “essential” missing element

• Represents a significant “vacuum” of knowledge

• Its discovery has the highest scientific urgency

The Physics of Hadron-Hadron Collisions

High energy hadron-hadron collisions

• Produce the highest-energy interactions

• Allow us to study qq̄′, gg and qg collisions

Have yielded some of the greatest insights

• Discovery of the W and Z bosons (UA1 & UA2)

• Discovery of the top quark (CDF & D0 / )

• Some of the most precise B physics measurements

• Precision electroweak measurements, eg. M_{W} and Γ_{W}

However, experiments are hard to perform

• Interaction cross section is large – ∼ 50 mb

• “Interesting” physics cross sections are small – σ_{b} ∼ 30 μb – σ_{W} ∼ 20 nb – σ_{t} ∼ 5 pb

• Events are complex

History of the CDF Experiment

CDF was initially approved in 1979!

• “Small” effort involving 16 institutions

• Had “engineering run” in October 1985

Followed by three successive runs

-I. Collected 30 nb^{−1} in 1987

0. Collected 4 pb^{−1} in 1988-89 I. Collected 100 pb^{−1} in 1992-95 (Run IA/IB)

– Made most precise M_{W} measurements

– Discovered the top quark

– Made first precise B physics measurements – Searched for host of new phenomena

Now starting Run II

• Initial goal is to collect 2 fb^{−1}

• Have significantly upgraded CDF ⇒ CDF II • Ultimate goal is to collect 10 − 15 fb^{−1}

Tevatron Upgrades for Run II

Tevatron has been significantly improved for Run II

• Main Injector replaces Tevatron for injection – Better p̄ transfer efficiency

– Emittance better-matched at injection

• Antiproton Recycler – effective doubling in p̄ production rate • Increase in √s from 1.8 to 2.0 TeV

⇒ 30-40% increase in production σ

• Go from 6 × 6 → 36 × 36 operation – Reduces interactions per crossing

Performance goals for collider are

• Increase typical L_{max} from 1.6 × 10^{31} to 2 × 10^{32} cm^{−2}s^{−1} • Deliver ∼ 2 fb^{−1} by 2002 (Run IIa)

• If successful in increasing p̄ availability – Go to 108 × 108 operation

– Increase L_{max} to 5 × 10^{32} cm^{−2}s^{−1}

– Deliver ∼ 15 fb^{−1} by end of 2007 (Run IIb)

CDF II Detector Upgrades

CDF has undergone major upgrades for Run II

• Tile-scintillating fibre calorimeters in 1.0 < |η| < 3.6 • Central outer tracker (COT)

• SVX II and ISL silicon tracking systems

• Fully pipelined electronics and new L1/L2/L3 triggers • Upgraded muon systems in 0.8 ≤ |η| ≤ 1.2

• C++ offline reconstruction and new processing farm

Canadian Contributions to CDF II

SVX II Mounting and Alignment

• Have provided SVX II barrel “adjustable” mounts – Assemble 3 barrels aligned to ∼ 100 μR

• “Inchworm” motors to locate SVX/ISL tracker parallel to beamline

– Use 5 pizeo-actuated motors making ∼ 10 μ steps – Keep 120 kg tracker parallel to beam

Calorimetry Software Management and Development

• P. Savard manages ALL calorimetry software development

• Responsible for – All calorimetry reconstruction

– Jet algorithm development/maintenance

– Missing E_{T} algorithms

– GFLASH calorimeter simulation

CDF II Physics Goals

Have a broad physics agenda

• Electroweak physics measurements, eg. – M_{W} to a precision of 20 MeV/c^{2}

– W W and W Z production

• QCD Physics – Extend studies of qq interactions up to √ŝ ∼ 1 TeV

• Exotics Searches – Extend SUSY searches to ∼ 300 GeV/c^{2}

– New vector bosons up to masses > 900 GeV/c^{2}

• Heavy Quark Physics – Precision top quark measurements

– Competitive B physics measurements

• Higgs Search

IPP focus is on top and bottom studies

• Some speculative Higgs studies

The State of Truth

CDF has now detected top in virtually all decay modes

• Lepton + Jets mode (for mass analysis) – lepton P_{T} > 20 GeV/c and E̸_{T} > 20 GeV

– ≥ 3 jets, E_{T} > 15 GeV and |η| < 2

– Require 1 b-tagged jet OR a 4th jet with E_{T} > 15 ⇒ 76 candidates (expect 31 ± 7 bkg)

• Dileptons – 2 leptons with P_{T} > 20 GeV/c

– Remove Drell-Yan and Z^{∘} background ⇒ 9 candidates (2.4 ± 0.5 events bkg)

• All hadronic decays – between 5 and 8 jets (E_{T} > 15 GeV and |η| < 2.0)

– ∑ E_{T} > 300 GeV + additional kinematic cuts ⇒ Observe 187 b-tagged jets (142 ± 11 expected bkg)

What are the Interesting Top Properties?

• Top Quark Mass – SM constrains it to M_{W} /M_{Z}

– Important consistency check of theory

• Top Quark Production Cross Section – Predicted by QCD with less uncertainties than σ_{b} – Look for single top production

• Top Quark Decay Rates – Measured t → W^{+} b/t → W^{+} q

– Look for t → H^{+} b

• Top Quark Decay Kinematics – SM expects decay to be polarized

– Have made first measurement of polarization

• Single Top Production

Top Quark Mass

Measurements now limited by statistical uncertainties

• Overall systematic uncertainty is ∼ 5 GeV/c^{2}

– Dominated by uncertainties in soft gluon contributions

and energy scale

• Run I statistics limit systematic uncertainty

The final Run I values are

M_{top} = 176.1 ± 6.6 GeV c^{2} (CDF)
M_{top} = 172.1 ± 7.1 GeV c^{2} (D0 / )
M_{top} = 174.3 ± 5.1 GeV c^{2} (RunI)

Latest W boson mass results

M_{W} = 80.433 ± 0.079 GeV c^{2} CDF
M_{W} = 80.482 ± 0.091 GeV c^{2} D0
M_{W} = 80.427 ± 0.046 GeV c^{2} LEP2

Top Quark Production Cross Section

The cross section obtained by a fit to

• Lepton + jets, dileptons and all hadronic event rates

• Depends on M_{top} because of acceptance

• Recent improvement is more accurate b-tagging efficiency

σ_{t t} = 6.5_{−1.4}^{+1.7} pb

• Theory predictions range from 4.7 to 5.5 pb

Top Physics Goals in Run II

Mass analysis uncertainties scale with statistics

• Can expect δM_{t} ∼ 3.0˜ GeV/c^{2}

• With δM_{W} ∼ 40 MeV/c^{2} – Begins to hint at the Higgs boson mass – Less than ∼ 210 GeV/c^{2}?

\frac{δM_{Higgs}}{M_{Higgs}} ∼ 40%

• Powerful “pointer” for Higgs searches

Measure angular correlations and decay kinematics

• Tell us about the electroweak force

• Polarization of W^{+} will be measurable

• Spin correlations (expect 90%) also detectable

Single top production

• Measures the “width” of the top quark

• Measure σ_{t} with 10% uncertainty

• Measure V_{tb} with 12% uncertainty

Expected Run IIa Result

CAP Congress, June 2001

Bottom Physics in Run II

Hadron-hadron collisions provide huge source of b quarks

• Will produce about 60 billion in Run IIA

• Challenge is to separate from enormous backgrounds

Run I has shown how this can be done

• Use “unique” signatures – Inclusive semileptonic decays to electrons or muons – Exclusive hadronic decays to J/ψ final states

∗ B → J/ψK/K^{∗}

∗ B_{s} → J/ψϕ

First to collect large sample of B^{0} → J/ψK_{s}^{0} decays

• Observed approximately 400 decays

sin(2β) = 0.79_{−0.44}^{+0.41}

CAP Congress, June 2001

CP Studies at Run II

Expect to make significant contributions to

• B lifetime measurements

• Spectroscopy of bq̄ systems

• CP violation is of most interest

Have developed B tagging techniques in Run I

• Jet charge tagging

• Soft lepton tagging

• Same-side tagging using “bachelor” π’s

• Run I “effective” tagging efficiency is

ϵD^{2} = 0.063 ± 0.017

• This will increase in Run II – Larger SVX II acceptance

– Lower lepton momentum thresholds – ϵD^{2} ∼ 0.091

• Results in δ(sin(2β)) ∼ 0.072

CP Violation and Mixing in B_{s} System

CDF II & D0 / have access to B_{s}^{0} system

• Measure CP violation in B_{s}^{0} → K^{+} K^{−} – Expect ϵD^{2} ∼ 0.11

– Will detect about 10,000 decays

– Gives a meaningful measurement of γ

Can perform detailed measurements of B_{s}^{0}

• Lifetime of B_{s}^{0} → J/ψϕ and B_{s}^{0} → D_{s}^{−}l^{+} ν_{l}X – Measures different CP states

– Yields 0.052 uncertainty on ΔΓ_{s}/Γ_{s}

– SM expectation is ∼ 0.15.

• Measure mixing in B_{s}^{0} → D_{s}^{−}π^{+} – Expect to make 5σ measurement of x_{s} out to 50-60.

Run II has comprehensive B physics programme

• Complementary to B Factory programme

• Has opportunity to compete....

Overview of Higgs Phenomenology

Higgs allows “spontaneous symmetry breaking” in SM

• Sets the EWK mass scale

• Yukawa interactions with fermions generates mass

It couples to quarks, leptons and EWK gauge bosons

• Production rate set by mass & couplings to “beam”

• Decays are mediated by same couplings

• In p p collisions, produced

– via qq̄ annihilations through ∗ virtual W/Z for light quark initial states

∗ directly via bb̄ initial states

– indirectly in gg interactions through gg → t t → H

Features of a Light SM Higgs

Higgs couplings to fermions & EWK bosons specify BR

• In region M_{H} < 130 GeV/c^{2}, H → bb̄ dominates

• For M_{H} > 130 GeV/c^{2}, H → W W^{∗} takes over

• Cross over is ∼ 135 GeV/c^{2}

• Other decay modes important, but are not relevant to a Tevatron search

Dicus, Steltzer, Sullivan & Willenbrock, PRD D59 (1999) Spira, private comm.

Higgs Production Cross Section

Cross section at Tevatron energies miniscule

• Arises from small couplings to light fermions • Makes “associated” production quite important

Some relevant points of reference:

• Total cross section ∼ 7 × 10^{10} pb

• Smallest measured total cross section ∼ 5 pb • Lowest upper limits ∼ 0.1 − 0.2 pb (W′)

Higgs Detection during Run II

SM Higgs production is in principle detectable

• Search of Run I data is clear demonstration – see G. Watts’ parallel talk

• For case of M_{H} = 120 GeV/c^{2}, cross section – for gg → H_{SM} is ∼ 0.9 pb

– for qq̄′ → W H_{SM} is ∼ 0.2 pb

– for qq̄ → Z H_{SM} is ∼ 0.1 pb

• Issues for detection are:

– Reconstruction technique and signal efficiency – Background rejection

Run II SUSY/Higgs Workshop has focussed activity

• Summarizes multiple studies done on – SM Higgs production

– MSSM Higgs scenarios

• Results provided in terms of ∫ Ldt

• Report from workshop now available (hep-ph/0010338)

B Tagging Issues

“B Tagging” an amalgam of various techniques

• Secondary vertex tagging primary technique (SECVTX) – Identify tracks with large impact parameter

– Search for 2 or 3-track combinations with large L_{xy}

• Probability of tracks coming from primary vertex (JPB tag)

• Identify soft leptons in core of b-jet (SLT)

Higgs Mass Resolution

Higgs mass resolution relies on jet reconstruction

• “Standard” CDF jet algorithm – gives M_{H} resolution ∼ 16 GeV/c^{2} at M_{H} = 110 GeV/c^{2}

• Track & shower-max data provides ∼ 20−30% improvement – Measured using photon-jet balancing

Assumed Improvement in Mass Resolution

Provides 10 − 20% improvement in H → bb̄ mass

This still requires work

• Topology dependence needs to be understood – Overlap with other objects in event – Combinatorial confusion

• Take into account other calorimetry effects

• Overall, promising development for Run II

Detection Strategy for Light SM Higgs

The most promising channels for a light SM Higgs are

p p → W^{+} H → l^{+} ν_{l} bb̄
p p → Z^{0} H → l^{+} l^{−} bb̄

• Identify e^{+} or μ^{+} with p_{T} > 20 GeV/c

• Require E̸_{T} > 20 GeV and reject dilepton events

• Only two jets with E_{T} > 25 GeV and > 10 GeV with |η| < 2 – Both b-tagged (one tight and one loose)

– Impose M_{bb} mass window ∼ ±15%

– No extra jets with E_{T} > 20 GeV and |η| < 2.4

Results in ϵ × B(H → bb̄) ∼ 2%

• Expected 8 fb observed for M_{H} = 90 GeV/c^{2} – Falls to 2 fb for M_{H} = 130 GeV/c^{2}

Background Rates for W H

All of the above come down to control of backgrounds

• Non-physical backgrounds estimated to be small

• Largest physical backgrounds are – QCD production of W bb̄

– Production of t t where two jets are missed

– Di-vector boson production

– Single top production (tb and tqb)

• Use HERWIG and PYTHIA, checked against Run I data Observed Cross Section (fb)

Process  M_{H} = 90  M_{H} = 110  M_{H} = 130
Signal  8.4  5.0  2.2
W bb̄  21.5  18.9  15.4
t t  9.6  12.2  13.0
V V  7.7  4.9  0.6
tb and tqb  9.4  11.8  12.7

Results for ZH Channel

ZH → l^{+} l^{−} bb̄ is straightforward

• Suffer about ×5 hit due to BR and lepton ID efficiency • Expect about 1/6 of W H observed rate

• Background rates are also approximately 1/6-1/7

ZH → ν_{l}ν̄_{l} bb̄ presents challenge

• Only have E̸_{T} signature from Z → νν̄

• Have two b-tagged jets reconstructing to H

• Creates a very “modest” signature – Issues with triggering

– Issues with beating down the backgrounds

Backgrounds to νν̄ bb̄ Final State

Physical backgrounds reasonably well understood

• QCD production of Zbb̄

• QCD production of W bb̄ (miss charged lepton)

• W Z and ZZ

• t t and single top

Observed cross section (fb)

Process  M_{H} = 90  M_{H} = 110  M_{H} = 130
Signal  3.6  2.7  0.9
Zbb̄  6.5  4.9  3.8
ZZ  5.4  3.4  0.0
t t  1.6  2.2  2.2
qq̄ → bb̄  ∼ 7  ∼ 6  ∼ 3

QCD bb̄ background difficult to estimate

• Depends crucially on E̸_{T} measurement

• Raw production cross section not well understood ⇒ Have to measure with data

Higher Mass Higgs Searches

Heavier Higgs (M_{H} > ∼ 130 GeV/c^{2}) rates lower

• H → W W^{∗} becomes dominant decay mode

• Production via gg → HX and qq̄′ → W H are important • Realistically, will require in excess of 15 fb^{−1}

This presents a number of decay channels

p p → H → W^{∗}W^{∗} → ll̄ν_{l}ν̄_{l}
p p → W^{+} H → W^{+} W^{∗}W^{∗} → l^{+} l^{+} qq̄′X

• These two represent the highest rate topologies

• Same-sign dileptons very clean

Signal and Background Estimates

Signal requirements for H → ll̄′ν_{l}ν̄_{l} are:

• Isolated leptons with p_{T} > 10 GeV/c and |η| < 1.5 • Require M_{ll′} < 78 GeV/c^{2} for e^{+} e^{−} and μ^{+} μ^{−}

– and M_{ll′} < 110 GeV/c^{2} for e^{+} μ^{−}

• Make decay angle cuts to reject τ^{+} τ^{−} events

• Reject events with high energy jets (eg., E_{T} > 95)

A more sophisticated likelihood analysis defines M_{C} ≡ √p_{T}^{2} (ll) + M^{2}(ll) + E̸_{T}

Heavy Higgs Signal and Background Rates

Estimates of signal and background rates

• Dominated by W^{+} W^{−} and t t

Observed Cross Section (fb)

Process  M_{H} = 140  M_{H} = 160  M_{H} = 180
Signal  2.6  1.5  1.0
Backgrounds  4.4  4.4  3.8
S/√B  2.1  3.9  2.7

Can estimate total integrated luminosity needed

• For either 3σ observation or 95% CL exclusion

• Include 10% systematic uncertainty

Associated Production of t tH

Recently, t tH has been reexamined at the Tevatron

Goldstein et al, hep-ph/0006311

• Analogous to bb̄H channel

• The t t gives efficient trigger and tag

Signal cross sections are small

• Estimate σ × B of t tH few fb

t tH Rates

Have to deal with very large t tX background

• t t + jj has rate of ∼ 1, 000 fb

• B-tagging is crucial to pull out signal – Require 2 jets, at least one b-tagged

• For M_{H} = 120 GeV/c^{2} and 15 fb^{−1} – 2-3 signal events and little background

Could yield a few “spectacular” events

Summary for SM Higgs Search

Results of the Workshop analysis can be combined:

• Use results of NN analysis

• Combine statistical power of both experiments

• Bands reflect 30% total additional “shift”

Summary

Run II will be the energy frontier for next 5-6 years

• Push luminosity and energy “envelope”

• Initial goal is 2 fb^{−1}, followed by ∼ 10 fb^{−1}

• State-of-the art detector

Collaboration positioned to push physics agenda • Plan for early results in spring/summer of 2002

• Aggressive team organized for assault on data

Will be exposing new features of nature

• Precision electroweak measurements

• Search for new phenomena at high mass scales

• New top quark measurements

• Contribute to B physics studies

• Press on the Higgs search